OpenStudy (anonymous):

2x^2 -9x+12≤ 3 I need help on this one

OpenStudy (anonymous):

subtract 3 from right

OpenStudy (anonymous):

Alright This is what I have

OpenStudy (anonymous):

x=3/2

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

[3/2,3]

OpenStudy (anonymous):

\[\frac{ 3 }{ 2 }\le x \le3\]

OpenStudy (anonymous):

So the solution would be written like [3/2, 2] ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Perfect! Thank you

OpenStudy (anonymous):

Will you be able to help me out with another q ?

OpenStudy (anonymous):

whats up?